# Benjamin S. Manning MIT John J. Horton

### v2

@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Horton, 2023; Argyle et al., 2022; Aher et al., 2023; Mei et al., 2024; Wang et al., 2025; Park et al., 2024, 2023; Chang et al., 2024; Binz and Schulz, 2023, 2024; Cerina and Duch, 2025; Manning et al., 2024; Capra et al., 2024; Shah et al., 2025; Anthis et al., 2025; Tranchero et al., 2024; Hansen et al., 2024; Broska et al., 2025) Across dozens of experiments, samples of these “AI subjects” respond with remarkable similarity to humans—even when simulating novel experiments that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2024; Li et al., 2024; Suh et al., 2025). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gao et al., 2024; Gui and Toubia, 2023) This inconsistency poses a challenge for AI simulations as robust and credible predictive models. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
A natural starting point is improving the instructions given to agents. These instructions, or “prompts,” are written descriptions given to the LLM specifying who it is, what it believes, or how it should behave and reason. Such second-person instructions (e.g., “You respond as a type-X person”) can profoundly affect output distributions because advanced LLMs have been explicitly fine-tuned to follow instructions (Bai et al., 2022; Ouyang et al., 2022). With an appropriate prompt (which can now be massive), many highly capable models can perform complex reasoning and mathematical tasks at levels sometimes better even than that of highly skilled humans. [^3] To date, Google’s Gemini 1.5 can accommodate 10 million tokens (Gemini Team, 2024). This is roughly equivalent to 15,000 pages, or about 20 copies of the Handbook of Experimental Economics (Kagel and Roth, 1995).
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025; Xie et al., 2025). As in other machine-learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a persona grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new environments. If the goal is to predict behavior in settings with no prior human data, how should researchers construct and evaluate prompts?
@@
In this paper, we build agents whose behavior in simulations usefully generalizes to what we see from humans across entire domains. Our approach mirrors what researchers generally try to do in social science, but in reverse. Rather than testing a theory with empirical data, theory is embedded in agents (via natural language instructions), which are then used to generate candidate data. The theory and agent composition is then optimized to reduce error with respect to real-world data from a domain where predictions are desired. Human data from distinct but conceptually similar settings serve as held-out test sets or are incorporated into training to improve generalization. We show that “general” agents constructed and validated in this way can improve the predictive power of AI subjects in novel settings. Both steps are essential: without theoretical grounding, optimized prompts may fail to meaningfully improve even in-sample predictions, and without cross-setting validation, they are prone to overfit.
The approach uses three kinds of data:
  1. 1.
    Training data (existing): Human data used to optimize AI simulations.
  2. 2.
    Validation or test data (existing): Human data generated from a distinct but presumably similar data-generation process as the training data. It is used to evaluate optimized agents.
  3. 3.
    Target data (novel): New human data not in the LLM’s training corpus, where we want to make predictions—from a presumably similar domain as the training and validation data.
@@
The second step is to optimize over this set to best match the human training data—there is an active literature on optimizing prompts (Khattab et al., 2024). We employ two optimization methods: (i) a selection method that identifies the optimal mixture of prompts from the candidate set (Xie et al., 2025; Leng et al., 2024; Bui et al., 2025), and (ii) a construction method that optimizes numerical parameters embedded directly in the prompts. Both methods are akin to the tuning of a machine learning model where the model’s parameters are optimized to best predict training data—or in this case, identify prompts that collectively best predict the human training data. This step produces a final set of prompts that we have reason to believe will generalize, because they achieve good in-sample fit while remaining grounded in theory. Poor fit suggests that the chosen theory does not adequately capture the relevant behavior.
@@
We assess whether prompts generalize using a train-test split approach inspired by the principles of invariant risk minimization (Peters et al., 2016; Heinze-Deml et al., 2018; Arjovsky et al., 2020). The final set of prompts is validated using settings distinct from the training data, where we also expect the same theory relevant to the target data to hold (e.g., optimize on a \$20 dictator game, but test on a \$5 dictator game). To be clear, this means the validation set necessarily comes from a distinct data-generation process from that which produced the training data. By construction, prompts with strong testing performance are then those that capture generalizable relationships predictive of human behavior across contexts. Consequently, if the novel target setting is governed by the same theory or causal mechanism used to construct the optimized prompts (e.g., the target is a \$50 dictator game), we may gain confidence that they will better predict human responses in that setting.
@@
We illustrate this approach and provide evidence of its efficacy with training and validation data drawn from experiments in the behavioral economics literature applied to AI subjects. All simulations use GPT-4o with temperature set to 1, though the approach is agnostic to the choice of model and hyperparameters. [^4] Gpt-4o, among the most widely used LLMs, and the temperature setting are defaults for the software used to run our simulations (Horton and Horton, 2024). We first apply the selection method to Arad and Rubinstein (2012)’s 11-20 money request game, where participants request an amount and receive a bonus if they choose exactly one less than their opponent. We use only the original dataset from the paper, which contained fewer than 200 observations. This setting is appealing to study because optimal play depends not only on the focal agent’s capabilities, but also on their beliefs about others’ strategic reasoning. Endowing AI subjects with distinct prompts corresponding to varying degrees of strategic reasoning—the theoretical focus of Arad and Rubinstein—produces a mixture that closely matches the original human data.
@@
When we validate these optimized samples on distinct variants of the 11-20 money request game, they are better predictors of human responses than baseline AI subjects with no additional instructions. By contrast, scientifically meaningless, or “atheoretical” AI subjects derived from historical figures, pseudo-scientific Myers-Briggs personality types, and those instructed to select particular numbers can sometimes match human distributions in one variant of the game but fail to generalize across others.
@@
We next test the predictive power of the optimized agents in settings where no prior human data exist. To do so, we construct four new games and collect responses from samples of preregistered crowdsourced participants (Horton et al., 2011) on Prolific. These games are derived from the original 11-20 game (and its variants), but adapted to other numeric ranges (1-10 and 1-7). The optimized sample of theoretically-motivated prompts produces responses that predict the new human data far better than the off-the-shelf baseline. Prediction error is decreased by 53%-73% across the games. Furthermore, these simulations predict the results of the new experiments in some games better than the most plausibly relevant human data from Arad and Rubinstein; in one case, the KL divergence is halved. By contrast, the alignment of the atheoretical prompts—which failed validation—with the new human data is often similar to or worse than the baseline AI.
@@
The first step is to establish a clearly defined space of experimental settings—for example, variants of public goods games, dictator games, or other allocation games differentiated by instructions, parameters, or other structural features. This idea is similar to that in the literature on program evaluation and external validity, where population-level treatment effects can be estimated by randomizing over a defined set of possible samples (Hotz et al., 2005; Allcott, 2015). From this space, a random subset of settings is drawn and randomly assigned to human subjects, who provide responses. [^5] This setup is also akin to the integrative experiment designs of Almaatouq et al. (2024). By comparing these human responses with LLM-generated predictions, we can make externally valid estimates across the entire space.
A key distinction from the setup laid out in Hotz et al. and Allcott is that appropriate coverage does not require that the family of settings share a data-generation process or that the underlying samples of human subjects are from the same population. Indeed, the variance of estimates naturally reflects how closely the held-out settings resemble those sampled. If all scenarios are variants of a single setting—e.g., a dictator game with different monetary amounts—performance is tightly bounded; if they span disparate domains—e.g., many types of allocation games—estimates are likely less precise. Note that even if a setting inadvertently appears in the model’s pre-training data, the random sampling procedure still accounts for its contribution—such cases may merely reduce prediction error.
@@ 2 Building prompts that generalize
This section explains the approach for constructing AI subjects that better approximate human response distributions in new environments. We begin with an extended example of predicting behavior in a novel public goods game where no previous human data exists. We then discuss the importance of validating across distinct data-generation processes and grounding candidate prompts in relevant social science theories. The section concludes with a summary of the approach and a brief description of the optimization methods used to construct agents.
@@ 2.1 A motivating example
One might think this problem could be addressed by first randomly splitting the human sample of dictator game offers into training and testing sets. Then identifying the best-performing prompts on the training set, and validating their performance on the test set (Mullainathan and Spiess, 2017; Ludwig et al., 2024). Indeed, an LLM instructed “You randomly choose numbers between 6 and 9” could reasonably predict both training and testing sets for any split of the human data better than the baseline LLM with no persona. However, such an approach will almost certainly fail to generalize beyond the specific data-generation process that produced the dictator game offers. It has, in effect, still overfit. Rather than overfitting to the training data, it has overfit to the whole data-generation process.
@@ 2.2 Identifying generalizable behavioral relationships
Instead, we propose leveraging principles from invariant risk minimization (Arjovsky et al., 2020) to identify behavioral relationships that remain stable despite shifts in the data-generating process. Rather than splitting a single dataset, we deliberately choose training data from one data-generation process (or several related processes) and validate using data from a related but distinct process from those used for training. Prompts that consistently predict behavior across these distinct datasets likely capture generalizable, and possibly causal (Peters et al., 2016; Heinze-Deml et al., 2018), drivers of behavior. As a result, these validated prompts should be more effective in predicting human responses in novel but theoretically similar settings.
Returning to our motivating example, suppose we have additional human data from a \$5 dictator game, with observed offers {1,1,1,2,2,2}. Although stakes differ, both the \$20 and \$5 dictator games likely share a common decision-making process: individuals give slightly less than half the available amount, a reasonable balance more similar to what is observed in humans (Henrich et al., 2001). By using the \$20 dictator game as training data and the \$5 dictator game for testing, we are implicitly filtering for prompts based on their capacity to predict this proportional response. The earlier atheoretical prompt “You randomly choose numbers between 6 and 9,” still fits the \$20 game perfectly, but clearly fails validation on the \$5 game. By contrast, the theoretically-motivated prompt “You are self-interested but fair,” likely generalizes effectively across both settings, and most importantly, the novel public goods game.
This validation process becomes even more robust when multiple distinct but theoretically-related datasets are available. Imagine dictator-game responses from \$1, \$5, \$10, \$20, \$50, and \$100 games, all exhibiting offers slightly below half. Optimizing and validating across these multiple settings reduces the risk of overfitting. With each new training data-generation process, it is less likely that an arbitrary prompt will generalize in-sample. [^7] We do not offer empirical examples of optimizing over multiple training settings in the main text (only single settings in Section 3). Appendix A provides such analyses, including an additional preregistered experiment. Thus, if the underlying relationship governing these settings also holds for a novel target setting, the validated prompts should robustly predict behavior there as well.
@@ 2.3 Theory as guide towards generalizability
Ultimately, applying a prompt—or a set of prompts—to a new environment requires an unavoidable inductive leap. What we can do is try to make the leap explicit and interpretable. LLMs are extremely good at following explicit, well-defined instructions. Because each prompt is a set of these flexible natural language instructions, researchers can both evaluate performance and reasonably assess its relevance for a new setting. [^10] Such flexibility also makes AI agents less susceptible to the Lucas critique (Lucas, 1976)—if at all. Unlike traditional agent-based models or mathematical theories, which have hard-coded predetermined processes to generate predictions, LLMs can, as we will show, interpret new policies in context, reason through their implications, and adaptively formulate responses. This mirrors how economic theory is used more broadly with real humans: it is tested against data, observed where it succeeds or fails, and cautiously extrapolated to settings just beyond current evidence. When a new tariff is proposed, for example, the theory motivating the tariff is not known to be universally “correct” in a complex dynamic world ex ante—it is a disciplined guess, shaped by assumptions and previous evidence. Theory-guided prompts, validated across environments, bring the same kind of disciplined reasoning to LLM-based predictions of human behavior.

### v3

@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Horton, 2023; Argyle et al., 2022; Aher et al., 2023; Mei et al., 2024; Wang et al., 2025; Park et al., 2024, 2023; Chang et al., 2024; Binz and Schulz, 2023, 2024; Cerina and Duch, 2025; Manning et al., 2024; Capra et al., 2024; Shah et al., 2025; Anthis et al., 2025; Hansen et al., 2024; Broska et al., 2025; Bybee, 2025; Kim et al., 2024) Across dozens of experiments, samples of these agents respond with remarkable similarity to humans—even when simulating novel studies that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2024; Li et al., 2024; Suh et al., 2025; Tranchero et al., 2024). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gao et al., 2024; Gui and Toubia, 2023) This inconsistency poses a challenge for AI simulations as robust and credible predictive models—particularly in settings where no prior human data exists. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
A natural starting point is improving the instructions given to agents. These instructions, or “prompts,” are written descriptions given to the LLM specifying who it is, what it believes, or how it should behave and reason. Such second-person instructions (e.g., “You respond as a type-X person”) can profoundly affect output distributions because advanced LLMs have been explicitly fine-tuned to follow instructions (Bai et al., 2022; Ouyang et al., 2022). With an appropriate prompt (which can now be massive), many highly capable models can perform complex reasoning and mathematical tasks at levels sometimes better even than those of highly skilled humans. [^3] To date, Google’s Gemini 1.5 can accommodate 10 million tokens (Gemini Team, 2024). This is roughly equivalent to 15,000 pages, or about 20 copies of the Handbook of Experimental Economics (Kagel and Roth, 1995).
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025; Xie et al., 2025). As in other machine-learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a prompt grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new settings. If the goal is to predict behavior in settings with no prior human data, how should researchers construct prompts?
@@
In this paper, we build agents whose behavior in simulations usefully generalizes to what we see from humans across entire domains. Our approach mirrors what researchers generally try to do in social science, but in reverse. Rather than testing a theory with empirical data, theory is embedded in agents (via natural language instructions), which are then used to generate candidate data. The theory and agent composition is then optimized to reduce error with respect to real-world data from a domain where predictions are desired. Human data from distinct but conceptually similar settings serve as held-out test sets or are incorporated into training to improve generalization. We show that “general” agents constructed and validated in this way can improve the predictive power of AI agents in novel settings. Both steps are essential: without theoretical grounding, optimized prompts may fail to meaningfully improve even in-sample predictions, and without cross-setting validation, they are prone to overfit.
The approach uses two kinds of data: i) training data—existing human data used to optimize AI simulations, and ii) validation or test data—existing human data related to but distinct from the training data, used to test whether the optimized agents generalize. The ultimate goal is to produce agents that can more accurately predict human behavior in novel target settings where no prior human data exist, but that lie in the same broad domain as the training and validation settings.
@@
The second step is to optimize over this set to best match the human training data. [^4] There is an active literature on optimizing prompts (Khattab et al., 2024). All candidate prompts are put in simulations of the settings that produced the training data (e.g., a \$20 dictator game). Optimization effectively filters these candidate prompts down to a subset whose simulated responses are closest to the human training distribution. We employ two optimization methods: (i) a selection method that identifies the optimal mixture of prompts from the candidate set (Xie et al., 2025; Leng et al., 2024; Bui et al., 2025), and (ii) a construction method that optimizes numerical parameters embedded directly in the prompts. Both methods produce a final set of prompts. If the set achieves good in-sample fit, we have reason to believe they will generalize because they are also grounded in a relevant theory. Poor in-sample fit suggests a mismatch between the theory and the training setting or an inadequate operationalization of that theory. In either case, we revise the candidate prompts and re-optimize.
@@
Given strong in-sample fit, we assess whether the final set of prompts generalize using a train-test split approach inspired by the principles of invariant risk minimization (Peters et al., 2016; Heinze-Deml et al., 2018; Arjovsky et al., 2020). The final set of prompts is placed in simulations of the settings that produced the human validation data, where we also expect the underlying theory to hold (e.g., optimize on a \$20 dictator game, but test on a \$5 dictator game). To be clear, this means the validation set necessarily comes from a distinct data-generating process from the one that produced the training data. We then compare the predictions from these simulations to the relevant human distributions. By construction, sets of prompts with strong test performance—those that accurately predict the validation data—are then those that capture generalizable relationships predictive of human behavior across contexts. Those that fail validation likely do not. Consequently, if the novel target setting is governed by the same theory or causal mechanism used to construct the optimized prompts (e.g., the target is a \$50 dictator game), we may gain confidence that they will better predict human responses in that setting.
@@
We illustrate this approach and provide evidence of its efficacy with training and validation data drawn from experiments in the behavioral economics literature applied to AI agents. All simulations use GPT-4o with temperature set to 1, though the approach is agnostic to the choice of model and hyperparameters. [^5] Gpt-4o, among the most widely used LLMs, and the temperature setting are defaults for the software used to run our simulations (Horton and Horton, 2024). We first apply the selection method to Arad and Rubinstein (2012)’s 11-20 money request game, where participants request an amount and receive a bonus if they choose exactly one less than their opponent. We use only the original dataset from the paper, which contained fewer than 200 observations. This setting is appealing to study because optimal play depends not only on the focal agent’s capabilities, but also on their beliefs about others’ strategic reasoning. Endowing AI agents with distinct prompts corresponding to varying degrees of strategic reasoning (the theoretical focus of Arad and Rubinstein) produces a mixture that closely matches the original human data.
@@
When we validate these optimized samples on distinct variants of the 11-20 money request game, they are better predictors of human responses than baseline AI agents with no additional instructions. By contrast, scientifically meaningless or “atheoretical” AI agents derived from historical figures, pseudo-scientific Myers-Briggs personality types, and those instructed to select particular numbers can sometimes match human distributions in one variant of the game but fail to generalize across others.
@@
We next test the predictive power of the optimized agents in target settings where no prior human data exist. To do so, we construct four new games and collect responses from samples of preregistered crowdsourced participants (Horton et al., 2011) on Prolific. These games are derived from the original 11-20 game (and its variants), but adapted to other numeric ranges (1-10 and 1-7). The optimized sample of theory-grounded prompts produces responses that predict the new human data far better than the off-the-shelf baseline. Prediction error is decreased by 53%-73% across the games. Furthermore, these simulations predict the results of the new experiments in some games better than the most plausibly relevant human data from Arad and Rubinstein; in one case, the KL divergence is halved. By contrast, the alignment of the atheoretical prompts—which failed validation—with the new human data is often similar to or worse than the baseline AI.
@@
The setup is very similar to the theoretical framework of Andrews et al. (2025), who study how well predictions from different economic and machine learning models transfer across economic domains. The first step is to establish a clearly defined space of experimental settings—for example, variants of public goods games, dictator games, or other allocation games differentiated by instructions, parameters, or other structural features. This idea is related to that in the literature on program evaluation and external validity, where population-level treatment effects can be estimated by randomizing over a defined set of possible samples (Hotz et al., 2005; Allcott, 2015; Dehejia et al., 2019). From this space, a random subset of settings is drawn and randomly assigned to human subjects, who provide responses. [^6] This is also similar to the integrative experiment designs of Almaatouq et al. (2024). By comparing these human responses with LLM-generated predictions, we can make externally valid estimates across the entire space.
A key distinction from the setups laid out in Hotz et al., Allcott, and Dehejia et al. is that appropriate coverage does not require that the family of settings share a data-generating process or that the underlying samples of human subjects are from the same population. Indeed, the variance of estimates naturally reflects how closely the held-out settings resemble those sampled. If all scenarios are variants of a single setting—e.g., a dictator game with different monetary amounts—performance is tightly bounded; if they span disparate domains—e.g., many types of allocation games—estimates are likely less precise. Note that even if a setting inadvertently appears in the model’s pre-training data, the random sampling procedure still accounts for its contribution—such cases may merely reduce prediction error.
@@ 2 Building prompts that generalize
This section explains the approach for constructing AI agents that better approximate human response distributions in new environments. We begin with an extended example of predicting behavior in a novel public goods game where no previous human data exists. We then discuss the importance of validating across distinct data-generating processes and grounding candidate prompts in relevant social science theories. The section concludes with a summary of the approach and a brief description of the optimization methods used to construct agents.
@@ 2.1 A motivating example
One might think this problem could be addressed by first randomly splitting the human sample of dictator game offers into training and testing sets. Then identifying the best-performing prompts on the training set, and validating their performance on the test set (Mullainathan and Spiess, 2017; Ludwig et al., 2024). Indeed, an LLM instructed “You randomly choose numbers between 6 and 9” could reasonably predict both training and testing sets for any split of the human data better than the baseline LLM with no persona. However, such an approach will almost certainly fail to generalize beyond the specific data-generating process that produced the dictator game offers. It has, in effect, still overfit. Rather than overfitting to the training data, it has overfit to the whole data-generating process.
@@ 2.2 Identifying generalizable behavioral relationships
Instead, we propose leveraging principles from invariant risk minimization (Arjovsky et al., 2020) to identify behavioral relationships that remain stable despite shifts in the data-generating process. Rather than splitting a single dataset, we deliberately choose training data from one data-generating process (or several related processes) and validate using data from a related but distinct process from those used for training. Prompts that consistently predict behavior across these distinct datasets likely capture generalizable, and possibly causal (Peters et al., 2016; Heinze-Deml et al., 2018), drivers of behavior. As a result, these validated prompts should be more effective in predicting human responses in novel but theoretically similar settings.
Returning to our motivating example, suppose we have additional human data from a \$5 dictator game, with observed offers {1,1,1,2,2,2}. Although stakes differ, both the \$20 and \$5 dictator games likely share a common decision-making process: individuals give slightly less than half the available amount, a reasonable balance more similar to what is observed in humans (Henrich et al., 2001). By using the \$20 dictator game as training data and the \$5 dictator game for testing, we are implicitly filtering for prompts based on their capacity to predict this proportional response. The earlier atheoretical prompt “You randomly choose numbers between 6 and 9,” still fits the \$20 game perfectly, but clearly fails validation on the \$5 game. By contrast, the theory-grounded prompt “You are self-interested but fair,” likely generalizes effectively across both settings, and most importantly, the novel public goods game.
This validation process becomes even more robust when multiple distinct but theoretically-related datasets are available. Imagine dictator-game responses from \$1, \$5, \$10, \$20, \$50, and \$100 games, all exhibiting offers slightly below half. Optimizing and validating across these multiple settings reduces the risk of overfitting. With each new training data-generating process, it is less likely that an arbitrary prompt will generalize in-sample. [^8] We do not offer empirical examples of optimizing over multiple training settings in the main text (only single settings in Section 3). Appendix A provides such analyses, including an additional preregistered experiment. Thus, if the underlying relationship governing these settings also holds for a novel target setting, the validated prompts should robustly predict behavior there as well.
@@ 2.3 Theory as guide towards generalizability
Ultimately, applying a prompt—or a set of prompts—to a new environment requires an unavoidable inductive leap. What we can do is try to make the leap explicit and interpretable. LLMs are extremely good at following explicit, well-defined instructions. Because each prompt is a set of these flexible natural language instructions, researchers can both evaluate performance and reasonably assess its relevance for a new setting. [^11] Such flexibility also makes AI agents less susceptible to the Lucas critique (Lucas, 1976)—if at all. Unlike traditional agent-based models or mathematical theories, which have hard-coded predetermined processes to generate predictions, LLMs can, as we will show, interpret new policies in context, reason through their implications, and adaptively formulate responses. This mirrors how economic theory is used more broadly with real humans: it is tested against data, observed where it succeeds or fails, and cautiously extrapolated to settings just beyond current evidence. When a new tariff is proposed, for example, the theory motivating the tariff is not known to be universally “correct” in a complex dynamic world ex ante. It is a disciplined guess, shaped by assumptions and previous evidence, which the economist then uses to make predictions about the world. Theory-guided prompts, validated across environments, bring the same kind of disciplined reasoning to simulated AI agent predictions of human behavior.